Circular No. 8897

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 17P/HOLMES H. H. Hsieh, A. Fitzsimmons, and D. L. Pollacco, Queen’s University,

Belfast, on behalf of the WASP Consortium, report that the SuperWASP-N facility was serendipitously imaging the position of comet 17P throughout October, and that the comet is not visible in unfiltered CCD images to a limiting magnitude of V ≈ 15 at or prior to Oct. 23.27 UT. The comet is visible in the next image taken on Oct. 23.99 and brightens steadily, becoming saturated at ≈ Oct. 24.10. From photometry performed using apertures 70′′ in radius (5.1 pixels; 124000 km at the comet), a nearly 3fold increase in the comet’s brightness is found (from V ≈ 9.7 to 8.6) in the 2.6 hours of unsaturated data. The brightening of the comet during this period is consistent with a power law with an exponent of 2, which would be expected from the linear expansion of an optically thick dust coma. Assuming a constant rate of expansion, the time of the comet’s intiial outburst is found to be ≈ Oct. 23.8 UT.

COMET P/2007 T4 (GIBBS) Additional astrometry and the following revised orbital elements from

MPEC 2007-V101 show this comet (cf. IAUC 8880) to be of short period:

T = 2007 July 20.3414 TT  ω = 42.^{o}3664
e = 0.614168  Ω = 37.1654 }2000.0
q = 2.005990 AU  i = 23.8690
a = 5.199127 AU n^{o} = 0.0831397 P = 11.9 years

COMET C/2007 T1 (McNAUGHT) Revised orbital elements (cf. IAUC 8877) from MPEC 2007-V53:

T = 2007 Dec. 12.520 TT  ω = 233.^{o}728
Ω = 111.429 }2000.0
q = 0.96882 AU  i = 117.646

2007 VW_{266}

Another retrograde asteroidal object (i = 108^{o}, q = 3.34 AU, P = 12.1 yr, mag ∼ 20) has been discovered, this time via the Mt. Lemmon survey (details given on MPEC 2007-W21).

2007 November 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green